OpenStudy (anonymous):

Multiply and simplify if possible. 4^√11 * 4^√10 a. 11 ^4√10 b. 11 c. 10 d. 4^√110

OpenStudy (zehanz):

You could use the following rule: \[\sqrt[n]{a} \cdot \sqrt[n]{b}=\sqrt[n]{ab}\]

OpenStudy (zehanz):

It means: the product of two alike n-th roots is the n-th root of the product, so\[\sqrt[4]{11} \cdot \sqrt[4]{10}=\sqrt[4]{11 \cdot 10}=...\] I'm pretty sure, you can see the right answer now!